# Injecting Measurement Information Yields a Fast and Noise-Robust Diffusion-Based Inverse Problem Solver

### v1

@@ 1 Introduction
with noise $\bm{\eta}$ and measurement operator $\mathcal{A}$. Known as inverse problems, such formulations appear in a multitude of fields, with applications including acoustic reconstruction (Kac 1966), seismic profiling (Hardage 1985), X-ray computed tomography and magnetic resonance imaging (Suetens 2017), and a large number of computer vision reconstruction tasks such as inpainting, deconvolution, colorization, super-resolution, and phase retrieval (Andrews and Hunt 1977).
@@
Often, inverting $\mathcal{A}$ is numerically intractable (Appendix C), meaning that solutions $\mathbf{x}$ satisfying $\mathcal{A}(\mathbf{x})=\mathbf{y}$ are not directly obtainable or unique (Vogel 2002). Moreover, due to measurement noise, it is often possible, but not practically desirable to fit perfectly to $\mathbf{y}$ for risk of overfitting to $\bm{\eta}$ (Aster, Borchers, and Thurber 2018). Therefore, a fundamental step in solving inverse problems is deciding how to select the best option from an equivalence class of solutions, i.e., choosing $\mathbf{x}_{*}\in\{\mathbf{x}:\mathcal{A}(\mathbf{x})\approx\mathbf{y}\}$.
In classical solvers, this is carried out by a regularizer on a normed error loss (Engl, Hanke, and Neubauer 1996):
@@
where $\epsilon$ is a soft error margin and $R$ is a simple function that satisfies user-specified heuristics, e.g., smoothness or total variation (Beck and Teboulle 2009). However, such approaches often fail to produce realistic results, as it can be difficult to design $R$ without introducing strong biases. With the advent of deep generative models, practitioners found that restricting solutions to the range of a generative model $G$ can greatly improve realism: let $\mathbf{x}=G(\mathbf{w})$ and optimize over $\mathbf{w}$, which can be latent inputs (Bora et al. 2017) or weights (Ulyanov, Vedaldi, and Lempitsky 2018) of a deep neural network. Overall, these methods improve the fidelity of $\mathbf{x}$, but they lack interpretability and require a judiciously selected $R$ and $\epsilon$.
@@
Recently, great strides have been made in solving inverse problems with diffusion models (Ho, Jain, and Abbeel 2020), which produce diverse, realistic samples (Dhariwal and Nichol 2021; Esser et al. 2024) with robust generalization guarantees (Kadkhodaie et al. 2023). Moreover, they are interpretable, directly modeling the (Stein) score $\nabla\log p_{t}$. Unconditional sampling proceeds by reversing a noising process on $\mathbf{x}_{0}\sim p_{\text{data}}$. Solvers then employ a conditional sampling process via a guidance term that pushes samples toward solutions consistent with $\mathbf{y}$:
@@
This approach faces a fundamental challenge: the guidance term depends on a consistency error $\mathcal{A}(\mathbf{x})-\mathbf{y}||$ that is only tractable for $\mathbf{x}=\mathbf{x}_{0}$ (Chung et al. 2022a). Such methods rely (explicitly or implicitly via (Song, Meng, and Ermon 2020)) on Tweedie’s formula, which estimates $\mathbf{x}_{0}$ given a noise prediction $\bm{\epsilon}_{t}\approx-\sigma_{t}\nabla\log p_{t}(\mathbf{x}_{t})$:
@@
This can then be substituted for $\mathbf{x}$ in the consistency error, producing a differentiable function with respect to $\mathbf{x}_{t}$.
However, a naive implementation of Eq. 3 introduces significant approximation error, as $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\mathbf{x}_{0}$ if and only if $\mathbf{x}_{t}$ is marginally normally distributed (Theorem 2.1). Existing methods use the unconditional score $\nabla\log p_{t}(\mathbf{x}_{t})$ where this assumption does not generally hold. In this work, we instead propose to use the conditional posterior mean $\mathbb{E}\left[\mathbf{x}_{0}|\mathbf{x}_{t},\,\mathbf{y}\right]$ as an estimate for $\mathbf{x}_{0}$ which requires estimating the forward process score $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$. While the forward process score is intractable during the reverse process, as $\mathbf{x}_{0}$ is unknown, we utilize the information in $\mathbf{y}$ in a statistically sufficient manner to obtain an estimate (Section 3).
@@
  • We demonstrate how this score can be plugged into any standard sampler (e.g., DDPM), resulting in an algorithm that is simple, noise-robust, neural backpropagation-free, and stable across time steps. Moreover, it achieves improvements in performance on a large selection of inverse problems, datasets and noise levels.  [^1] Code for method and experiments provided in https://anonymous.4open.science/r/diffusion˙conditional˙sampling.
@@ 2 Revisiting Tweedie’s for Inverse Problems
Diffusion models (Ho, Jain, and Abbeel 2020) reverse a noise-corrupting forward process with marginals $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$,
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\underbrace{\sqrt{1-\alpha_{t}}}_{\sigma_{t}}\mathbf{z},\hskip 18.06749pt\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (4)
parameterized by a monotonically time-decreasing scalar $\alpha_{t}$. New samples are generated via the reverse diffusion process which leverages the learned score function $s_{\theta}=-\sigma_{t}^{-1}\bm{\epsilon}_{\theta}\approx\nabla\log p_{t}(\mathbf{x}_{t})$ (Anderson 1982; Vincent 2011; Song et al. 2020). Diffusion-based solvers for inverse problems can be categorized as posterior or projection solvers, and aim to modify the reverse process such that the final variate $\mathbf{x}_{0}$ coincides with a member of the solution set $\{\mathbf{x}_{0}:\mathcal{A}(\mathbf{x}_{0})\approx\mathbf{y}\}$ [^2] We defer a more extensive discussion on diffusion models diffusion-based inverse problem solvers to Appendix A.. This paradigm is afflicted by a fundamental computability paradox: since the consistency error is only explicitly known at $t=0$ via the likelihood function
@@
we cannot exactly guide the diffusion process at time $t>0$ without first solving for $\mathbf{x}_{0}$. Simultaneously, we cannot generally obtain $\mathbf{x}_{0}$ without first computing $\mathbf{x}_{t}$. Accurately estimating $\hat{\mathbf{x}}_{0}\approx\mathbf{x}_{0}$ is a fundamental problem all solvers must contend with to function properly.
In posterior solvers, the computation of $\nabla\log p(\mathbf{y}|\mathbf{x}_{t})$, is approximated by $\nabla\log p(\mathbf{y}|\hat{\mathbf{x}}_{0})$. In projection solvers, this is hidden in the projection step $\mathbf{P}\mathbf{x}_{t}$, which is driven by a projection on $\hat{\mathbf{x}}_{0}$, followed by a DDIM step (Song, Meng, and Ermon 2020) that involves $\hat{\mathbf{x}}_{0}$. In both cases, Tweedie’s formula is used to create an estimate for $\mathbf{x}_{0}$, given the current $\mathbf{x}_{t}$ by predicting the posterior mean
@@
4:  $\bm{\epsilon}\sim\mathcal{N}(\textbf{0},\mathbf{I})$
@@
6:  $\mathbf{x}_{t-1}\leftarrow\texttt{ddpm\_step}(\mathbf{x}_{t},\epsilon_{\theta}+\epsilon_{y})$
@@
1:Input: $\mathbf{y},\mathcal{A},\mathbf{x}_{t},\epsilon_{\theta}$$|$ Output: $\epsilon_{y}$
@@
5:  $\epsilon_{y}\leftarrow\epsilon_{y}+\eta\nabla\log p_{t}\left(\mathbf{y}|\frac{\mathbf{x}_{t}-\sigma_{t}(\epsilon_{\theta}+\epsilon_{\mathbf{y}})}{\sqrt{\alpha_{t}}}\right)$
6:  $\hat{\mathbf{x}}\leftarrow\texttt{Tweedie's}(\mathbf{x}_{t},\epsilon_{\theta}+\epsilon_{y})$
@@ 3 Diffusion Conditional Sampling
We propose a novel framework for solving inverse problems with diffusion models via a measurement consistent Tweedie’s formula (Eq. 7). At each step, we form a single-parameter measurement model whose maximum likelihood estimator approximates the forward process score $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ (Section 3.1). This estimator is optimized with a noise-robust, likelihood-based early stopping criterion (Section 3.2). The learned score is then input to a standard DDPM sampler (Ho, Jain, and Abbeel 2020), resulting in Algorithm 1. This approach is motivated by both powerful theoretical guarantees (Section 3.3), and yields significant computational advantages (Section 3.4).
@@ 3.1 Measurement Likelihood Model
In standard unconditional sampling, this is also known as the forward process $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ (Ho, Jain, and Abbeel 2020). However, when running the reverse process (unconditional sampling) this is a non-trivial quantity to evaluate, as $\mathbf{x}_{0}$ is unknown. We apply Tweedie’s to Eq. 8 and solve for the forward process score via a closed form expression for the likelihood (Eq. 5):
@@ 3.2 Noise-Aware Maximization
Formally, we optimize Eq. 9 until the likelihood of the alternate hypothesis, $\mathbb{H}_{1}$ (that res is not distributed as $\eta$), is below a desired threshold $p_{\text{critical}}$. Since overfitting is more problematic at the end of sampling ($t\approx 0$) than the beginning of sampling ($t\approx T$), we set $p_{\text{critical}}$ dynamically as a function of $t$, namely $p_{\text{critical}}(t)=\sigma_{t}$. This scheme is heavily inspired by the classical two-sided $\mathbf{z}$-test (Hogg et al. 2013) with $d$ samples, where $d$ is the dimensionality of the image. We use the early-stopping criterion at each time $t$
@@
where $\xi_{i}\overset{\text{iid}}{\sim}\mathcal{N}(\mathbf{0},\sigma_{\mathbf{y}}^{2})$ and $\Phi$ is the CDF of a standard normal distribution. We note that this differs from a classical $\mathbf{z}$-test since we are not seeking to reject the null hypothesis, but optimizing until the null hypothesis can no longer be rejected with sufficiently high probability. The full noise-aware maximization algorithm can be summarized by Algorithm 2. Since our loss function (Eq. 9) is quadratic, our proposed nam has worst-case linear convergence guarantees due to classical results with gradient descent (Boyd and Vandenberghe 2004; Ryu and Boyd 2016) with linear $\mathcal{A}$.
@@ A Near-Linear Reverse Process
In Figure 8, we validate the robustness of our algorithm to the total diffusion steps ($T$) with the super-resolution task on a subset of the FFHQ $256\times 256$ dataset. We compare against DPS (Chung et al. 2023), DPS-JF (a neural backpropagation-free variant of DPS), and DDNM (Wang, Yu, and Zhang 2022) with additive Gaussian noise at standard deviation $\sigma_{\mathbf{y}}=0.05$.
@@
FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring
$\sigma_{\mathbf{y}}=0.01$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$
Ours 0.137 30.14 19.45 0.024 34.84 21.19 0.088 25.11 19.25 0.103 28.69 22.62 0.087 29.48 26.67
MCG 0.144 24.84 31.47 0.073 30.59 22.22 0.453 15.44 185.54 0.209 23.51 67.88 0.217 22.93 292.1
DDNM 0.208 26.28 51.33 0.040 33.08 23.35 0.209 18.12 88.32 0.235 26.09 71.47 0.424 14.22 250.9
LGD-MC 0.238 23.45 39.55 0.272 23.46 57.70 0.372 15.45 86.00 0.405 18.78 64.31 0.520 13.90 106.0
DPS 0.163 25.91 33.21 0.105 29.54 29.72 0.113 23.52 24.41 0.129 26.48 26.85 0.159 24.41 29.84
RED-Diff 0.178 29.81 45.68 0.035 33.72 25.03 0.090 25.20 19.98 0.234 29.72 52.09 0.191 29.14 116.9
FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring
$\sigma_{\mathbf{y}}=0.1$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$
Ours 0.175 24.88 30.11 0.149 27.54 32.80 0.163 23.22 26.44 0.176 25.96 26.08 0.224 24.61 31.40
MCG 0.546 20.44 102.6 0.227 26.00 50.40 0.579 15.30 207.2 0.429 25.80 69.29 0.973 -7.104 295.3
DDNM 0.623 21.49 145.9 0.179 24.96 39.18 0.334 19.20 72.11 1.220 10.73 176.8 0.739 5.099 524.0
LGD-MC 0.256 22.31 39.58 0.288 22.22 56.05 0.384 15.38 87.72 0.415 18.30 66.04 0.524 13.65 105.4
DPS 0.185 24.79 35.46 0.157 26.72 35.24 0.158 22.58 32.47 0.180 24.720 33.53 0.212 22.41 35.09
RED-Diff 0.665 22.10 155.1 0.155 28.62 34.78 0.298 22.96 61.14 0.447 26.93 106.3 0.423 24.16 120.1
ImageNet SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring
$\sigma_{\mathbf{y}}=0.01$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$
Ours 0.238 23.45 39.41 0.142 26.06 34.46 0.230 20.63 37.11 0.253 24.22 38.96 0.203 24.62 38.63
MCG 0.638 15.62 89.39 0.198 24.34 35.19 0.273 16.68 80.35 0.645 21.18 124.6 0.980 -5.726 231.1
DDNM 0.333 25.16 51.33 0.084 28.35 20.27 0.258 17.42 85.41 0.456 24.35 67.98 0.694 5.721 304.2
LGD-MC 0.662 14.460 113.5 0.650 14.57 129.8 0.696 11.63 133.9 0.796 10.46 165.7 0.807 9.609 184.7
DPS 0.309 23.99 49.81 0.266 25.05 38.87 0.301 18.76 34.85 0.493 19.14 61.59 0.460 18.65 53.21
RED-Diff 0.386 25.07 57.06 0.090 28.17 16.71 0.239 19.99 54.38 0.459 24.70 68.71 0.376 23.66 55.77
ImageNet SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring
$\sigma_{\mathbf{y}}=0.1$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$
Ours 0.402 22.99 48.21 0.166 26.04 34.47 0.243 19.70 46.03 0.407 22.28 51.13 0.435 20.43 61.48
MCG 0.886 14.01 145.1 0.459 19.92 78.86 0.433 15.63 124.0 0.650 22.00 117.4 0.984 -6.868 231.3
DDNM 0.751 20.98 133.3 0.1693 25.63 35.72 0.400 18.06 110.8 1.221 9.602 202.7 0.783 5.009 350.1
LGD-MC 0.671 14.01 116.51 0.661 14.19 131.2 0.701 11.53 134.58 0.804 10.44 167.3 0.806 9.587 185.4
DPS 0.540 18.63 85.06 0.506 20.10 82.74 0.479 18.03 83.06 0.412 20.57 65.07 0.450 18.91 75.65
RED-Diff 0.747 20.66 136.35 0.167 25.38 32.99 0.374 19.68 88.20 0.660 23.19 110.9 0.591 21.27 138.8

### v2

@@ 1 Introduction
with noise $\bm{\eta}$ and measurement operator $\mathcal{A}$. Known as inverse problems, such formulations appear in a multitude of fields, with applications including acoustic reconstruction (Kac,, 1966), seismic profiling (Hardage,, 1985), X-ray computed tomography and magnetic resonance imaging (Suetens,, 2017), and a large number of computer vision reconstruction tasks such as inpainting, deconvolution, colorization, super-resolution, and phase retrieval (Andrews and Hunt,, 1977).
@@
Often, inverting $\mathcal{A}$ is numerically intractable (Appendix C), meaning that solutions $\mathbf{x}$ satisfying $\mathcal{A}(\mathbf{x})=\mathbf{y}$ are not directly obtainable or unique (Vogel,, 2002). Moreover, due to measurement noise, it is often possible, but not practically desirable to fit perfectly to $\mathbf{y}$ for risk of overfitting to $\bm{\eta}$ (Aster et al.,, 2018). Therefore, a fundamental step in solving inverse problems is deciding how to select the best option from an equivalence class of solutions, i.e., choosing $\mathbf{x}_{*}\in\{\mathbf{x}:\mathcal{A}(\mathbf{x})\approx\mathbf{y}\}$.
In classical solvers, this is carried out by a regularizer on a normed error loss (Engl et al.,, 1996):
@@
where $\epsilon$ is a soft error margin and $R$ is a simple function that satisfies user-specified heuristics, e.g., smoothness or total variation (Beck and Teboulle,, 2009). However, such approaches often fail to produce realistic results, as it can be difficult to design $R$ without introducing strong biases. With the advent of deep generative models, practitioners found that restricting solutions to the range of a generative model $G$ can greatly improve realism: let $\mathbf{x}=G(\mathbf{w})$ and optimize over $\mathbf{w}$, which can be latent inputs (Bora et al.,, 2017) or weights (Ulyanov et al.,, 2018) of a deep neural network. Overall, these methods improve the fidelity of $\mathbf{x}$, but they lack interpretability and require a judiciously selected $R$ and $\epsilon$.
@@
Recently, great strides have been made in solving inverse problems with diffusion models (Ho et al.,, 2020), which produce diverse, realistic samples (Dhariwal and Nichol,, 2021; Esser et al.,, 2024) with robust generalization guarantees (Kadkhodaie et al.,, 2023). Moreover, they are interpretable, directly modeling the (Stein) score $\nabla\log p_{t}$. Unconditional sampling proceeds by reversing a noising process on $\mathbf{x}_{0}\sim p_{\text{data}}$. Solvers then employ a conditional sampling process via a guidance term that pushes samples toward solutions consistent with $\mathbf{y}$:
@@
This approach faces a fundamental challenge: the guidance term depends on a consistency error $\mathcal{A}(\mathbf{x})-\mathbf{y}||$ that is only tractable for $\mathbf{x}=\mathbf{x}_{0}$ (Chung et al., 2022a, ). Such methods rely (explicitly or implicitly via (Song et al., 2020a, )) on Tweedie’s formula, which estimates $\mathbf{x}_{0}$ given a noise prediction $\bm{\epsilon}_{t}\approx-\sigma_{t}\nabla\log p_{t}(\mathbf{x}_{t})$:
@@
This approximation is then substituted for $\mathbf{x}$ in the consistency error, producing a differentiable function with respect to $\mathbf{x}_{t}$.
This naive implementation of Eq. 3 introduces significant approximation error, as $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\mathbf{x}_{0}$ if and only if the marginal distribution of $\mathbf{x}_{t}$ is normally distributed (Theorem 2.1). Existing methods use the unconditional score $\nabla\log p_{t}(\mathbf{x}_{t})$ where this assumption does not generally hold. In this work, we instead propose to use the conditional posterior mean $\mathbb{E}\left[\mathbf{x}_{0}|\mathbf{x}_{t},\,\mathbf{y}\right]$ as an estimate for $\mathbf{x}_{0}$ which requires estimating the forward process score, $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$. While the forward process score is intractable during the reverse process (inference), since $\mathbf{x}_{0}$ is unknown, we can utilize the information contained in $\mathbf{y}$ in a statistically sufficient manner to obtain an estimate (Section 3).
@@
  • We demonstrate how this score can be plugged into any standard sampler (e.g., DDPM), resulting in an algorithm that is simple, noise-robust, neural backpropagation-free, and stable across time steps. Moreover, it achieves improvements in performance on a large selection of inverse problems, datasets and noise levels  [^1] Code for method and experiments in the link..
@@ 2 Revisiting Tweedie’s for Inverse Problems
Diffusion models Ho et al., (2020) reverse a noise-corrupting forward process, with marginals $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$, defined as
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\underbrace{\sqrt{1-\alpha_{t}}}_{\sigma_{t}}\mathbf{z},\hskip 7.22743pt\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (4)
and parameterized by a monotonically time-decreasing scalar $\alpha_{t}$. New samples are generated via the reverse diffusion process which leverages the learned score function $s_{\theta}=-\sigma_{t}^{-1}\bm{\epsilon}_{\theta}\approx\nabla\log p_{t}(\mathbf{x}_{t})$ (Anderson,, 1982; Vincent,, 2011; Song et al., 2020b, ). Diffusion-based solvers for inverse problems can be categorized as posterior or projection solvers, and aim to modify the reverse process such that the final variate $\mathbf{x}_{0}$ coincides with an element of the solution set $\{\mathbf{x}_{0}:\mathcal{A}(\mathbf{x}_{0})\approx\mathbf{y}\}$ [^2] We defer a more extensive discussion on diffusion models diffusion-based inverse problem solvers to Appendix A.. This paradigm is afflicted by a fundamental computability paradox: since the consistency error is only explicitly known at $t=0$ by applying the likelihood function
@@
we cannot exactly guide the diffusion process at time $t>0$ without first solving for $\mathbf{x}_{0}$. Simultaneously, we cannot generally obtain $\mathbf{x}_{0}$ without first computing $\mathbf{x}_{t}$. Accurately estimating $\hat{\mathbf{x}}_{0}\approx\mathbf{x}_{0}$ is a fundamental challenge both types of solvers must contend with to function properly.
In posterior solvers, the computation of $\nabla\log p(\mathbf{y}|\mathbf{x}_{t})$, is approximated by $\nabla\log p(\mathbf{y}|\hat{\mathbf{x}}_{0})$. In projection solvers, this approximation is hidden in the projection step $\mathbf{P}\mathbf{x}_{t}$, which is driven by a projection on $\hat{\mathbf{x}}_{0}$, followed by a DDIM step (Song et al., 2020a, ) that involves $\hat{\mathbf{x}}_{0}$. In both cases, Tweedie’s formula is used to create an estimate for $\mathbf{x}_{0}$, given the current $\mathbf{x}_{t}$ by predicting the posterior mean
@@
4:  $\bm{\epsilon}\leftarrow\epsilon_{\theta}(\mathbf{x}_{t})$
@@
6:  $\mathbf{x}_{t-1}\leftarrow\texttt{ddpm\_step}(\mathbf{x}_{t},\bm{\epsilon}+\epsilon_{\mathbf{y}})$
@@
1:Input: $\mathbf{y},\mathcal{A},\mathbf{x}_{t},\bm{\epsilon}$$|$ Output: $\epsilon_{y}$
@@
5:  $\epsilon_{y}\leftarrow\epsilon_{y}+\eta\nabla\log p_{t}\left(\mathbf{y}|\frac{\mathbf{x}_{t}-\sigma_{t}(\bm{\epsilon}+\epsilon_{\mathbf{y}})}{\sqrt{\alpha_{t}}}\right)$
6:  $\hat{\mathbf{x}}\leftarrow\texttt{Tweedie's}(\mathbf{x}_{t},\bm{\epsilon}+\epsilon_{y})$
@@ 3 Diffusion Conditional Sampling
We propose a novel framework for solving inverse problems with diffusion models via a measurement consistent Tweedie’s formula (Eq. 7). At each step, we form a single-parameter measurement model whose maximum likelihood estimator approximates the forward process score $\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ (Section 3.1). This estimator is optimized with a noise-robust, likelihood-based early stopping criterion (Section 3.2). The learned score is then input to a standard DDPM sampler (Ho et al.,, 2020), resulting in Algorithm 1. This approach is motivated by both powerful theoretical guarantees (Section 3.3), and yields significant computational advantages (Section 3.4).
@@ 3.1 Measurement Likelihood Model
In standard unconditional sampling, this is also known as the forward process $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ (Ho et al.,, 2020). However, when running the reverse process (unconditional sampling) this is a non-trivial quantity to evaluate, as $\mathbf{x}_{0}$ is unknown. We apply Tweedie’s to Eq. 8 and solve for the forward process score via a closed form expression for the likelihood (Eq. 5):
@@ 3.2 Noise-Aware Maximization
Formally, we optimize Eq. 9 until the likelihood of the alternate hypothesis, $\mathbb{H}_{1}$ (that res is not distributed as $\eta$), is below a desired threshold $p_{\text{critical}}$. Since overfitting is more problematic at the end of sampling ($t\approx 0$) than the beginning of sampling ($t\approx T$), we set $p_{\text{critical}}$ dynamically as a function of $t$, namely $p_{\text{critical}}(t)=\sigma_{t}$. This scheme is heavily inspired by the classical two-sided $\mathbf{z}$-test (Hogg et al.,, 2013) with $d$ samples, where $d$ is the dimensionality of the image. We use the early-stopping criterion at each time $t$
@@
where $\xi_{i}\overset{\text{iid}}{\sim}\mathcal{N}(\mathbf{0},\sigma_{\mathbf{y}}^{2})$ and $\Phi$ is the CDF of a standard normal distribution. We note that this differs from a classical $\mathbf{z}$-test since we are not seeking to reject the null hypothesis, but optimizing until the null hypothesis can no longer be rejected with sufficiently high probability. The full noise-aware maximization algorithm can be summarized by Algorithm 2. Since our loss function (Eq. 9) is quadratic, our proposed nam has worst-case linear convergence guarantees due to classical results with gradient descent (Boyd and Vandenberghe,, 2004; Ryu and Boyd,, 2016) with linear $\mathcal{A}$.
@@ A Near-Linear Reverse Process
In Figure 8, we validate the robustness of our algorithm to the total diffusion steps ($T$) with the super-resolution task on a subset of the FFHQ $256\times 256$ dataset. We compare against DPS Chung et al., (2023), DPS-JF (a neural backpropagation-free variant of DPS), and DDNM Wang et al., (2022) with additive Gaussian noise at standard deviation $\sigma_{\mathbf{y}}=0.05$.
@@
FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring $\sigma_{\mathbf{y}}=0.01$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.137 30.14 19.45 0.024 34.84 21.19 0.088 25.11 19.25 0.103 28.69 22.62 0.087 29.48 26.67 MCG 0.144 24.84 31.47 0.073 30.59 22.22 0.453 15.44 185.54 0.209 23.51 67.88 0.217 22.93 292.1 DDNM 0.208 26.28 51.33 0.040 33.08 23.35 0.209 18.12 88.32 0.235 26.09 71.47 0.424 14.22 250.9 LGD-MC 0.238 23.45 39.55 0.272 23.46 57.70 0.372 15.45 86.00 0.405 18.78 64.31 0.520 13.90 106.0 DPS 0.163 25.91 33.21 0.105 29.54 29.72 0.113 23.52 24.41 0.129 26.48 26.85 0.159 24.41 29.84 RED-Diff 0.178 29.81 45.68 0.035 33.72 25.03 0.090 25.20 19.98 0.234 29.72 52.09 0.191 29.14 116.9 FFHQ SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring $\sigma_{\mathbf{y}}=0.1$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.175 24.88 30.11 0.149 27.54 32.80 0.163 23.22 26.44 0.176 25.96 26.08 0.224 24.61 31.40 MCG 0.546 20.44 102.6 0.227 26.00 50.40 0.579 15.30 207.2 0.429 25.80 69.29 0.973 -7.104 295.3 DDNM 0.623 21.49 145.9 0.179 24.96 39.18 0.334 19.20 72.11 1.220 10.73 176.8 0.739 5.099 524.0 LGD-MC 0.256 22.31 39.58 0.288 22.22 56.05 0.384 15.38 87.72 0.415 18.30 66.04 0.524 13.65 105.4 DPS 0.185 24.79 35.46 0.157 26.72 35.24 0.158 22.58 32.47 0.180 24.720 33.53 0.212 22.41 35.09 RED-Diff 0.665 22.10 155.1 0.155 28.62 34.78 0.298 22.96 61.14 0.447 26.93 106.3 0.423 24.16 120.1 ImageNet SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring $\sigma_{\mathbf{y}}=0.01$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.238 23.45 39.41 0.142 26.06 34.46 0.230 20.63 37.11 0.253 24.22 38.96 0.203 24.62 38.63 MCG 0.638 15.62 89.39 0.198 24.34 35.19 0.273 16.68 80.35 0.645 21.18 124.6 0.980 -5.726 231.1 DDNM 0.333 25.16 51.33 0.084 28.35 20.27 0.258 17.42 85.41 0.456 24.35 67.98 0.694 5.721 304.2 LGD-MC 0.662 14.460 113.5 0.650 14.57 129.8 0.696 11.63 133.9 0.796 10.46 165.7 0.807 9.609 184.7 DPS 0.309 23.99 49.81 0.266 25.05 38.87 0.301 18.76 34.85 0.493 19.14 61.59 0.460 18.65 53.21 RED-Diff 0.386 25.07 57.06 0.090 28.17 16.71 0.239 19.99 54.38 0.459 24.70 68.71 0.376 23.66 55.77 ImageNet SR $\times 4$ Random Inpainting Box Inpainting Gaussian Deblurring Motion Deblurring $\sigma_{\mathbf{y}}=0.1$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ Ours 0.402 22.99 48.21 0.166 26.04 34.47 0.243 19.70 46.03 0.407 22.28 51.13 0.435 20.43 61.48 MCG 0.886 14.01 145.1 0.459 19.92 78.86 0.433 15.63 124.0 0.650 22.00 117.4 0.984 -6.868 231.3 DDNM 0.751 20.98 133.3 0.1693 25.63 35.72 0.400 18.06 110.8 1.221 9.602 202.7 0.783 5.009 350.1 LGD-MC 0.671 14.01 116.51 0.661 14.19 131.2 0.701 11.53 134.58 0.804 10.44 167.3 0.806 9.587 185.4 DPS 0.540 18.63 85.06 0.506 20.10 82.74 0.479 18.03 83.06 0.412 20.57 65.07 0.450 18.91 75.65 RED-Diff 0.747 20.66 136.35 0.167 25.38 32.99 0.374 19.68 88.20 0.660 23.19 110.9 0.591 21.27 138.8